OpenStudy (anonymous):

I am really having a difficult time with my online math class right now, so if you could explain what I'm supposed to do I'd really appreciate it. Thank you! Find the 35th term of the sequence in which a1 = -10 and the common difference is 4.

jimthompson5910 (jim_thompson5910):

an = a1 + d(n-1) an = -10 + 4(n-1) ... plug in the given info a35 = -10 + 4(35-1) ... now plug in n = 35 since you want the 35th term a35 = -10 + 4(34) a35 = -10 + 136 a35 = 126 So the 35th term is 126

OpenStudy (anonymous):

Thank you so much! I kept looking in my lesson reading but I couldn't find it. You are awesome.

jimthompson5910 (jim_thompson5910):

you're welcome, glad it makes sense now